OpenStudy (anonymous):

simplify the expression (4x^2/6y^4)^3

OpenStudy (lgbasallote):

\[\LARGE (\frac{4x^2}{6y^4})^3 \implies \frac{(4x^2)^2}{(6y^4)^3} \implies \frac{4^2 x^{2 \times 2}}{6^3 y^{4\times 3}}\] lol i hope you understand that

OpenStudy (lgbasallote):

oops lol made a mistake

OpenStudy (lgbasallote):

let me fix it

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

\[\huge (\frac{4x^2}{6y^4})^3 \implies \frac{(4x^2)^3}{(6y^4)^3} \implies \frac{4^3 x^{2 \times 3}}{6^3 y^{4 \times 3}}\] i hope you understand that lol

OpenStudy (anonymous):

kind of thanks

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

(4x^2/6y^4)3 = (64x^6/216x^12)=(8/216x^6)